# Beware of Calibration Data for Pruning Large Language Models

### v1

@@ 6.4 Whether self-generated synthetic calibration data is more similar to training data?
where , is the expectation and standard deviation of the next token’s log probability given the prefix , respectively. refers to choosing the bottom of subsequences based on scores from the sequence . Thus, the higher a sample’s Min-K%++ score, the more likely it is to appear in the training data. Figure 5 shows the Min-50%++ score distribution of C4, Wikipedia, Slimpajama and our self-generated synthetic data. We can clearly observe that the self-generated synthetic data has higher Min-50%++ scores than the other calibration data. It indicates that the self-generated synthetic calibration data is indeed similar to the training data, confirming the validity of using self-generated data as a proxy for the training data.

### v2

@@ 6.4 Whether self-generated synthetic calibration data is more similar to training data?
where , is the expectation and standard deviation of the next token’s log probability given the prefix , respectively. refers to choosing the bottom of subsequences based on scores from the sequence . Thus, the higher a sample’s Min-K%++ score, the more likely it is to appear in the training data. Figure 5 uses kernel density estimation to show the distribution of Min-K%++ values for C4, Wikipedia, SlimPajama and our self-generated synthetic data. We can clearly observe that the self-generated synthetic data has higher Min-50%++ scores than the other calibration data. It indicates that the self-generated synthetic calibration data is indeed similar to the training data, confirming the validity of using self-generated data as a proxy for the training data.
@@ 7 Conclusion and Future Work
Acknowledgments
We want to thank all the anonymous reviewers for their valuable comments. This work was supported by the National Science Foundation of China (NSFC No. 62206194), the Natural Science Foundation of Jiangsu Province, China (Grant No. BK20220488), the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), and the Priority Academic Program Development of Jiangsu Higher Education Institutions.
@@ Appendix B More Results of synthetic calibration data
Method | Data | Alpaca () | BoolQ | Winogrande | PIQA | Hellaswag | ARC-e | ARC-c | MMLU | Avg. | |
Wanda | Wiki | 10.42 | 66.80 | 63.84 | 70.55 | 56.69 | 64.78 | 34.23 | 22.94 | 54.26 | |
w/ Syn | 9.62 | 68.29 | 64.40 | 71.49 | 58.89 | 64.73 | 35.41 | 24.01 | 55.32 | +1.06 | |
C4 | 10.42 | 66.30 | 64.50 | 71.12 | 58.92 | 64.92 | 33.91 | 23.06 | 54.68 | ||
w/ Syn | 10.07 | 67.46 | 64.15 | 71.38 | 59.05 | 65.64 | 33.83 | 23.92 | 55.06 | +0.38 | |
Slim | 10.23 | 66.83 | 63.68 | 71.10 | 57.54 | 64.68 | 33.98 | 22.95 | 54.39 | ||
w/ Syn | 9.92 | 67.91 | 64.63 | 71.45 | 58.52 | 65.28 | 33.93 | 23.29 | 55.00 | +0.61 | |
DCLM | 9.88 | 68.92 | 64.25 | 71.15 | 58.72 | 64.81 | 33.98 | 23.65 | 55.07 | ||
w/ Syn | 9.77 | 68.90 | 64.56 | 71.71 | 58.90 | 65.24 | 34.47 | 23.61 | 55.34 | +0.27 | |
DSnoT | Wiki | 10.92 | 66.24 | 62.72 | 70.55 | 55.55 | 64.10 | 33.16 | 23.05 | 53.62 | |
w/ Syn | 10.40 | 65.44 | 64.01 | 71.49 | 57.77 | 64.86 | 34.30 | 23.90 | 54.54 | +0.92 | |
C4 | 10.88 | 65.25 | 64.04 | 71.22 | 57.15 | 64.40 | 32.82 | 23.45 | 54.05 | ||
w/ Syn | 9.90 | 66.18 | 64.72 | 71.00 | 57.19 | 64.86 | 33.45 | 24.87 | 54.61 | +0.56 | |
Slim | 10.76 | 65.66 | 63.66 | 70.82 | 56.17 | 64.43 | 32.51 | 23.15 | 53.77 | ||
w/ Syn | 10.04 | 65.23 | 63.22 | 70.84 | 56.56 | 65.11 | 33.11 | 23.67 | 53.97 | +0.20 | |
DCLM | 10.37 | 66.65 | 63.99 | 71.44 | 56.77 | 64.56 | 33.30 | 23.73 | 54.35 | ||
w/ Syn | 9.82 | 66.24 | 64.01 | 71.00 | 57.64 | 64.86 | 33.70 | 24.09 | 54.51 | +0.16 | |
OWL | Wiki | 9.30 | 66.50 | 66.05 | 71.82 | 61.90 | 67.57 | 35.89 | 26.07 | 56.54 | |
w/ Syn | 9.13 | 69.85 | 66.38 | 73.18 | 62.86 | 67.89 | 35.07 | 26.34 | 57.37 | +0.83 | |
C4 | 9.19 | 66.73 | 67.34 | 72.74 | 62.87 | 67.54 | 35.68 | 26.20 | 57.02 | ||
w/ Syn | 9.11 | 68.47 | 67.40 | 72.52 | 62.99 | 66.75 | 35.41 | 27.28 | 57.26 | +0.24 | |
Slim | 9.21 | 67.52 | 66.91 | 72.32 | 62.25 | 66.70 | 34.91 | 26.05 | 56.67 | ||
w/ Syn | 9.04 | 69.30 | 67.80 | 72.31 | 62.56 | 67.17 | 35.75 | 26.86 | 57.39 | +0.72 | |
DCLM | 9.08 | 69.79 | 67.94 | 72.39 | 62.73 | 67.06 | 35.85 | 26.45 | 57.46 | ||
w/ Syn | 9.10 | 69.57 | 67.72 | 72.63 | 62.60 | 67.59 | 35.84 | 26.32 | 57.47 | +0.01 |